AAVSOnet status, 2015-11-01

Sun, 11/01/2015 - 18:51

BSM_Berry, BSM_Hamren, BSM_HQ and BSM_South are fully operational.  After the polar alignment for BSM_HQ, all images are sharp and untrailed.

BSM_NM is waiting for mount repair parts.  These were shipped last week, so if they fix the mount problem, then BSM_NM may be back on the air shortly.  The New Mexico weather has been poor for the past couple of months, so very few imaging hours have been missed.

SRO and TMO61 are fully operational.  Gary Walker and Jon Holtzman just finished testing the new LISA spectrograph on TMO61, which will be available for shared risk projects during the next full moon.  Thanks to Bart Staels for proving the funding for this new instrument!

Coker30 is having declination axis problems right now, and because Bill Goff will be on travel off/on through the Fall meeting, may not be back on the air for a couple of weeks.

OC61 has been off-line recently to have its mirror realuminized.  It should begin to take data again this week.